OpenStudy (anonymous):

What is the domain of x/(4x+9)?

OpenStudy (abdullahm):

The domain is all real numbers except for the values of x that make the denominator equal to 0.

OpenStudy (anonymous):

Ok, so far I've only thought of -9/4. So (-inf, -9/4)u(-9/4, inf)

OpenStudy (abdullahm):

Absolutely perfect!

OpenStudy (anonymous):

Awesome, thank you!

OpenStudy (abdullahm):

\[(- \infty , -\frac{9}{4})\cup (-\frac{9}{4} , \infty)\]